OpenStudy (anonymous):

When combining like terms, how do I know whether to add or subtract, etc.? For example: (5x^2 - 7x - 7) + (7x^2 + 2x + 5) Do I add 5x^2 and 7x^2, as well as 7x & 2x? Please help. I'm overwhelmingly confused. x.x

hartnn (hartnn):

the sign with 5x^2 is + the sign in the middle of the brackets is also + the sign with 7x^2 is + so yes, you would add 5x^2+7x^2 to get (5+7)x^2. the sign with -7x is - the sign in the middle of the brackets is also + the sign with +2x is + so, you will do this : -7x+2x = (-7+2)x =... ?

hartnn (hartnn):

i only factored out 'x' from -7x+2x to get (-7+2)x. then there will be no confusion whether to add/subtract, just solve the bracket -7+2 =... ?

hartnn (hartnn):

did you understand ?

hartnn (hartnn):

by the way , \(\huge \color{red}{\text{Welcome to Open Study}}\ddot\smile\)

OpenStudy (anonymous):

yes, thank you a lot!